OpenStudy (anonymous):

The line 3x-2y=12 has a) a slope of 3/2 and a y intercept of -6 b) a slope of -3/2 and a y intercept of 6 c) a slope of 3 and a y intercept of -2 d) a slope of -3 and a y intercept of -6

OpenStudy (anonymous):

i think the answer is either a or b!

OpenStudy (anonymous):

why dont you try rearranging your equation such that you get it in the form y = mx + b

OpenStudy (anonymous):

ok lets see.....first you have to add -2y on both sides?

OpenStudy (anonymous):

no, you are trying to remove -2y from the left hand side. so you add +2y to both sides so that your LHS becomes 3x-2y+2y -----> 3x+0

OpenStudy (anonymous):

2y=3x-12 then divide them with 2 y=3/2 x -6 a

OpenStudy (anonymous):

y=mx+c c = y intercept m = slope y=3/2 x -6 y = m x + c clear? m=3/2 , c =-6